Gluten sensitivity affects approximately 15% of people in the us.You decide to randomly sample 800 individuals and have them report their gluten intolerance. 96 people report gluten sensitivity. 


1
Expert's answer
2021-05-14T12:02:42-0400

Let we have to calculate the probability that 115 individual in the sample with wheat intolerance.


Probability of wheat intolerance is p = 0.15

Sample size, n = 800


As np = 800"\times" 0.15 = 120 > 10 and n(1-p) = 800"\times" (1-0.15) = 680 > 10,


we can use the normal approximation to the binomial distribution with continuity correction.


a. Let X be the number of individuals in the sample with wheat intolerance.


X~N(np, np(1-p) or X~N(120, 102)


P(X = 115) = P(114.5 < x<115.5)


"=P(\dfrac{(114.5-120)}{\sqrt{102}}< z< \dfrac{(115.5-120)}{\sqrt{102}})"


"=P(-0.5446< z< -0.4456)"


= 0.293